OpenStudy (anonymous):

How its possible \(\frac{n-1}{n+1} = 1 - \frac{2}{n+1}\)

OpenStudy (blockcolder):

Your goal is to get another n+1 in the numerator so that it cancels out. \[\frac{n-1}{n+1}=\frac{n+1-2}{n+1}=\frac{n+1}{n+1}-\frac{2}{n+1}\]

OpenStudy (anonymous):

hmm i think i get it, thank you :)